OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (anonymous):

\[3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}=10\sqrt{2x}\\ \sqrt{2x}(3+2x-15-10)=0\\ \sqrt{2x}(2x-22)=0\\ x=0\qquad OR\qquad 2x-22=0\implies x=11 \]

OpenStudy (anonymous):

so is that the anwser or just part of it ?

OpenStudy (anonymous):

those are the steps to follow

OpenStudy (anonymous):

and answer is at the end

OpenStudy (anonymous):

ty